OpenStudy (anonymous):

the path traveled by a bottlenose dolphin as it jumps out of the water is modled by the equation y=0.4x2+3x, where y is the height above water and x is the horizontal distance in feet. If a beam of light is shown upward at an angle modeled by the equation x+y=10, at what height from the water's surface will the beam of light hit the dolphin

OpenStudy (aum):

y = 0.4x^2 + 3x x + y = 10 or y = -x + 10 To find where the two graphs will intersect, equate the two y's and solve for x. Then, put x in the second equation and find y. 0.4x^2 + 3x = -x + 10 solve for x first. Then find y.

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.